The average lifetime of smoke detectors that a company manufactures is 5 years, or 60 months, and the standard deviation is 6 months. Find the probability that a random sample of 35 smoke detectors will have a mean lifetime between 59 and 62 months. Assume that the sample is taken from a large population and the correction factor can be ignored. Round the final answer to four decimal places and intermediate z-value calculations to two decimal places P (59 < X < 62) =
Answer #1

µ =    60                              
σ =    6                              
n=   35                              
we need to calculate probability for ,                                  
59   ≤ X ≤    62                          
X1 =    59   ,   X2 =   62                  
                                  
Z1 =   (X1 - µ )/(σ/√n) =   -0.99                          
Z2 =   (X2 - µ )/(σ/√n) =   1.97                          
                                  
P (   59   < X <    62   ) =    P (    -0.99   < Z <    1.97   )
                                  
= P ( Z <    1.97   ) - P ( Z <   -0.99   ) =    0.9757   -    0.1621   =    0.8136